A valószínűségszámítás II. gyakorlaton való részvétel

dátum:  IX. 10. IX. 17. IX. 24. X. 1. X. 8. X. 15. X. 29.

diákok:

Bóka Zsolt  +  +  +  +  +  +  +
Csaba Anikó  +  +  +  +  +  +  −
Darvas Gábor  −  +  +  −  +  +  −
Fábián Andrea  +  +  +  +  +  +  +
Fazekas Zoltán  −  +  +  −  −  −  −
Filko Attila  −  −  −  −  +  −  −
Garai Krisztina  −  −  −  −  +  +  +
Gombos Gábor  +  −  −  +  +  +  +
Gorzsás József  +  −  −  −  −  −  −
Hadházi István  +  +  +  +  −  +  +
Héver Timea  +  −  −  +  +  +  +
Homoki István  +  −  −  +  +  +  −
Keresztes György  −  +  +  −  −  −  −
Kis–Tóth Attila  +  +  +  +  +  +  +
Kovács Marianna  +  −  −  +  +  +  +
László Anikó  +  −  −  −  −  −  −
Lukács Attila  +  +  +  +  +  −  +
Major Imre  −  +  +  +  +  +  −
Makra Tamás  −  −  −  +  +  +  +
Malbik Erzsébet  −  +  +  +  +  −  +
Mészáros Fruzsina  +  +  +  +  +  +  −
Nagy Erzsébet Bernadett  +  +  +  +  +  +  +
Nyerges Beáta  +  +  +  +  +  +  +
Oláh Erika  +  −  −  −  +  −  +
Orosz Erzsébet  +  −  −  +  +  +  +
Ottmár Ilona  +  +  +  +  +  +  +
Ölbei Kinga  +  +  +  +  −  +  −
Peleskey Eleonóra  +  +  +  +  +  +  +
Rab Konstantin  −  +  +  −  +  −  +
Szabó Attila  +  +  +  +  +  +  +
Szuetta Judit  +  +  +  +  +  +  +
Tarr Levente  −  +  +  −  −  −  −
Tóth Krisztina  −  −  +  −  −  +  −
Veligdán Szilvia  −  +  +  +  +  +  +
Vécsei Timea  −  +  +  −  −  +  −
Völfinger Réka  +  +  +  −  +  +  +
Zákány Irma Ildikó  −  +  +  +  +  +  +
Zsitnyár Judit  +  −  −  −  −  −  −

dátum:  X1. 5. XI. 12. XI. 19. XI. 26. XII. 10.

diákok:

Bóka Zsolt  −  +  +  +  +
Csaba Anikó  +  +  +  +  +
Darvas Gábor  +  −  +  −  −
Fábián Andrea  −  +  +  +  +
Garai Krisztina  −  +  +  −  +
Gombos Gábor  +  +  −  +  +
Hadházi István  −  +  +  +  +
Héver Timea  +  −  +  +  −
Homoki István  −  −  −  +  +
Kis–Tóth Attila  +  +  +  +  +
Kovács Marianna  +  −  +  +  −
Lukács Attila  +  −  +  +  +
Major Imre  −  +  +  +  +
Makra Tamás  +  +  +  +  +
Malbik Erzsébet  +  −  +  −  +
Mészáros Fruzsina  +  +  +  +  +
Nagy Erzsébet Bernadett  −  +  +  −  +
Nyerges Beáta  +  +  +  +  +
Oláh Erika  +  +  +  +  +
Orosz Erzsébet  +  −  +  +  +
Ottmár Ilona  −  −  +  −  +
Ölbei Kinga  +  +  −  +  +
Peleskey Eleonóra  −  −  +  +  +
Rab Konstantin  −  +  +  −  −
Szabó Attila  +  −  −  +  +
Szuetta Judit  +  +  +  +  +
Tóth Krisztina  −  +  +  +  −
Veligdán Szilvia  +  +  +  +  +
Vécsei Timea  −  −  −  −  −
Völfinger Réka  +  −  −  −  −
Zákány Irma Ildikó  +  +  +  +  +